dr = b(cos ϕdr + rsin ϕdϕ) note: b is a constant 


Separable DE


1
Expert's answer
2021-12-13T05:48:47-0500

"dr\ \ =\ \ b\left(\mathrm{cos}\left(\phi \right)dr\ \ \ +\ \ r\mathrm{sin}\left(\phi \right)d\phi \right) \\

 \\

\mathrm{1}\ \ =\ \ b\left(\mathrm{cos}\left(\phi \right)\ \ \ +\ \ r\mathrm{sin}\left(\phi \right)\frac{d\phi }{dr}\right) \\

 \\

\frac{\left(\mathrm{1}-b\mathrm{cos}\left(\phi \right)\right)}{b\ \mathrm{sin}\left(\phi \right)}\ \ =\ \ \left(\ r\frac{d\phi }{dr}\right) \\

 \\

b\frac{\mathrm{sin}\left(\phi \right)\ \ d\phi }{\left(\mathrm{1}-b\mathrm{cos}\left(\phi \right)\right)}\ \ =\ \ \left(\ \frac{dr}{r}\right) \\

 \\

\int{\frac{d\left(\mathrm{1}-b\mathrm{cos}\left(\phi \right)\ \right)\ }{\left(\mathrm{1}-b\mathrm{cos}\left(\phi \right)\right)}}\ \ =\ \ \int{d\left(\mathrm{ln}\left(r\right)\right)} \\

 \\

\mathrm{ln}\left(\mathrm{1}-b\mathrm{cos}\left(\phi \right)\ \right)\ \ \ =\ \ \mathrm{ln}\left(r\right)+\mathrm{ln}\left(C\right) \\

 \\

\mathrm{ln}\left(\mathrm{1}-b\mathrm{cos}\left(\phi \right)\ \right)\ \ \ =\ \ \mathrm{ln}\left(Cr\right) \\

 \\

\mathrm{1}-b\mathrm{cos}\left(\phi \right)\ =Cr\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \left(where\ \ C\ \ is\ \ an\ \ abitrarily\ \ \ cons\mathrm{tan}t\right)"